OpenStudy (anonymous):

The area of circle M is 16π, and the radius of circle N is twice the radius of circle M. What is the area of circle N? 8 8π 32π 64 64π

OpenStudy (ghazi):

pie* r^2= 16 * pie r=4 therefore radius of N= 2*4=8